# MaskGCT: Zero-Shot Text-to-Speech with Masked Generative Codec Transformer

### v1

@@ 4.2.1 Comparison with Baselines
We compare MaskGCT with baselines in terms of similarity, robustness, and generation quality. The main results are shown in Table 1. MaskGCT demonstrates excellent performance on all metrics and achieves human-level similarity, naturalness, and intelligibility. In similarity, MaskGCT’s SIM and SMOS both outperform the best baseline, whether assessed using the total length of ground truth or the predicted total duration (0.67$\rightarrow$0.687 in LibriSpeech, 0.643$\rightarrow$0.717 in SeedTTS test-en, 0.75$\rightarrow$0.774 in SeedTTS test-zh for SIM; +0.016 in LibriSpeech, +0.762 in SeedTTS test-en, +0.550 in SeedTTS test-zh for SMOS). When compared with human recordings, MaskGCT achieves human-level similarity across all three test sets (+0.017, -0.002, and +0.027 for SIM respectively in the three test sets, and +0.156, +0.319, and +0.244 for SMOS respectively in the three test sets). In robustness, MaskGCT likewise results nearly on par with ground truth (with 2.634, 2.623, 2.273 WER on LibriSpeech, SeedTTS test-en, and SeedTTS test-zh, respectively), exhibiting enhanced robustness compared to AR-based models and performing on par or better than NAR-based models such as VoiceBox and NaturalSpeech 3, without relying on phone-level duration predictions. In generation quality, MaskGCT achieves +0.109, +0.033, +0.050, and +0.083 CMOS across the three test sets when compared with human recordings, indicating that MaskGCT attains human-level naturalness on these test sets. We also observe that MaskGCT exhibits excellent performance when using both ground truth total duration and predicted total duration, indicating the robustness of MaskGCT within a reasonable range of total speech duration and the capability of our total duration predictor to yield appropriate durations.
4.2.2 Autoregressive vs. Masked Generative Models
We compare MaskGCT to replacing T2S MaskGCT with an AR T2S model (which we call AR + SoundStorm). Table 2 shows the performance of these two models on all three test sets. MaskGCT demonstrates improved similarity, robustness, and CMOS (+0.126 on LibriSpeech test-clean, +0.051 on SeedTTS test-en, and +0.323 on SeedTTS test-zh) across all three test sets. We also conducted comparisons on more challenging hard cases (such as repeating words, and tongue twisters, which are often considered as samples where TTS systems are prone to hallucinations). MaskGCT exhibited a more pronounced robustness advantage in these scenarios. See details in the Appendix K. In addition, compared to AR-based models, MaskGCT offers the capability to control the total duration of the generated speech, along with fewer inference steps, requiring only 25 to 50 steps for T2S models to achieve optimal results for speeches of any length. Conversely, the inference steps for AR-based models increase linearly with the length of the speech.
4.2.3 Duration Length Analysis
@@
AR + SoundStorm 0.672 3.267 0.998 4.203 -0.017
MaskGCT 0.687 2.634 0.886 4.274 0.109
@@
AR + SoundStorm 0.683 2.846 0.323 4.026 -0.051
MaskGCT 0.717 2.623 0.188 4.242 0.000
@@
AR + SoundStorm 0.747 3.865 0.238 3.775 -0.323
MaskGCT 0.774 2.273 0.106 4.094 0.000
@@
Figure 3: WER vs. Total Duration Multiplier (Speech Tempo) for SeedTTS test-en and SeedTTS test-zh test sets. We use inference timesteps as 25 for the T2S model.
@@ 4.3 Ablation Study
Inference Timesteps. Figure 4 shows the relationship between inference steps and metrics SIM and WER for SeedTTS test-zh (left) and test-en (right). Initially, SIM increases significantly, stabilizing after 25 steps. For test-zh, SIM rises from 0.761 at 5 steps to 0.771 at 75 steps, and for test-en, from 0.696 to 0.715. SIM reaches high values with just 10 steps but peaks around 25 steps. WER improves more dramatically, especially up to 25 steps. For test-zh, WER drops from 10.19 at 5 steps to 2.507 at 25 steps, and for test-en, from 8.096 to 2.346. Both SIM and WER show minimal changes beyond 25 steps. These findings indicate that while SIM metrics can be sufficiently optimized with around 10 inference steps, achieving the lowest WER values requires approximately 25 inference steps. Beyond this threshold, both SIM and WER metrics exhibit minimal changes, implying that further increases in inference steps do not yield substantial improvements in these performance metrics. Therefore, for practical applications, 25 inference steps may be considered optimal for balancing SIM and WER, ensuring efficient and effective performance.
@@
Text Tokenizer. We compare two different methods of text tokenization: (1) Grapheme-to-Phoneme(G2P), which converts text into a sequence of phonemes; (2) Byte Pair Encoding (BPE). See more details for the methods of text tokenization in Appendix D. The results indicate that G2P outperforms BPE in English with a higher similarity score (SIM) of 0.728 compared to 0.711 and a lower Word Error Rate (WER) of 2.466 versus 4.036. Conversely, in Chinese, G2P maintains a slightly higher similarity score (0.777 vs. 0.769) but BPE achieves a lower WER (1.921 vs. 2.338). These findings suggest that while G2P is superior in preserving text similarity and reducing errors in English, BPE is more effective in minimizing WER in Chinese. We hypothesize that the reason might be that the Chinese G2P system we used still has deficiencies in handling polyphonic characters, whereas BPE is capable of learning different pronunciations for the same character based on context.
Table 4: Text tokenizer: G2P vs. BPE. In this ablation study, we utilize the ground truth speech length.
SIM-O $\uparrow$ WER $\downarrow$
SeedTTS test-en
G2P 0.728 2.466
BPE 0.711 4.036
SeedTTS test-zh
G2P 0.777 2.183
BPE 0.769 1.921
@@ 5 Conclusion
In this paper, we present MaskGCT, a large-scale zero-shot TTS system that leverages fully non-autoregressive masked generative codec transformers while not requiring precise alignment information between text and speech. MaskGCT achieves high-quality text-to-speech synthesis using text to predict semantic tokens extracted from a speech self-supervised learning (SSL) model, and then predicting acoustic tokens conditioned on these semantic tokens. Our experiments demonstrate that MaskGCT outperforms the state-of-the-art TTS system on speech quality, similarity, and intelligibility with scaled model size and training data, and MaskGCT can control the total duration of generated speech. We also explore the scalability of MaskGCT in tasks such as speech translation, voice conversion, emotion control, and speech content editing, demonstrating the potential of MaskGCT as a foundational model for speech generation.
@@ Appendix A Details of MaskGCT
Model Architecture. We use a Llama-style [Touvron et al., 2023] Transformer architecture as the backbone of our model, incorporating gated linear units with GELU [Hendrycks and Gimpel, 2016] activation (SwiGLU), rotation position encoding [Su et al., 2024], etc., but replacing causal attention with bidirectional attention. We also use adaptive RMSNorm [Zhang and Sennrich, 2019], which accepts the time step $t$ as the condition. Table 5 presents the key hyperparameters of the models.
@@
Inference Steps for the S2A model. The S2A model generates tokens layer by layer during inference. Since the acoustic codec follows an RVQ structure, we can view the S2A inference as a process from coarse to fine. We also use more iterations in the initial layers, as the first few layers carry more information. By default, we use inference steps of $[40,16,1,1,1,1,1,1,1,1,1,1]$ for each layer, however, we find that the S2A model can also perform well with fewer steps, such as $[10,1,1,1,1,1,1,1,1,1,1,1]$, with only a very slight performance loss.
@@
Appendix B Details of Semantic/Acoustic Codec
For semantic codec, we train a VQVAE model using the hidden features from the 17th layer of W2v-BERT 2.0, incorporating factorized codec [Lezama et al., 2022] technology. The original hidden dimension of 1,024 is projected into a lower-dimensional space for quantization. The codebook size is set to 8,192, with a codebook dimension of 8. We employ only the $\mathcal{L}_{1}$ loss as the reconstruction target, optimizing the codebook with codebook loss and commitment loss. The input features are normalized to have a mean of 0 and a variance of 1, based on the statistics of the training dataset. The encoder and the decoder are each composed of 12 mirrored ConvNext blocks, featuring a kernel size of 7 and a hidden size of 384.
@@ Appendix B Details of Semantic/Acoustic Codec
For acoustic codec, the basic architecture of the encoder follows Kumar et al. [2024] and the decoder follows Siuzdak [2023]. The Vocos-based decoder can model amplitude and phase, enabling waveform generation through inverse STFT transformation without requiring upsampling. The number of RVQ layers, codebook size, and codebook dimension are set to 12, 8,192, and 8, respectively. We utilize the multi-scale mel-reconstruction loss [Kumar et al., 2024] $\mathcal{L}_{\text{rec}}$, for the adversarial loss $\mathcal{L}_{\text{adv}}$, we employ both the multi-period discriminator (MPD) and the multi-band multi-scale STFT discriminator, as proposed by Lee et al. [2022], Kumar et al. [2024]. Additionally, we incorporate the relative feature matching loss $\mathcal{L}_{\text{feat}}$. For codebook learning, we use the codebook loss $\mathcal{L}_{\text{codebook}}$ and the commitment loss $\mathcal{L}_{\text{commit}}$ from VQ-VAE. We set $\lambda_{\text{rec}}=10.0$, $\lambda_{\text{adv}}=2.0$, $\lambda_{\text{feat}}=2.0$, $\lambda_{\text{codebook}}=1.0$, $\lambda_{\text{commit}}=0.25$ as coefficients for balancing each loss terms. Figure 5 shows the overview of the semantic codec and acoustic codec, Table 7 presents the detailed model configurations of semantic codec and acoustic codec.
@@
Appendix C Details of Duration Predictor

### v2

@@ 4.2.1 Comparison with Baselines
We compare MaskGCT with baselines in terms of similarity, robustness, and generation quality. The main results are shown in Table 2. MaskGCT demonstrates excellent performance on all metrics and achieves human-level similarity, naturalness, and intelligibility. In similarity, MaskGCT’s SIM-O and SMOS both outperform the best baseline, whether assessed using the total length of ground truth or the predicted total duration (0.67$\rightarrow$0.687 in LibriSpeech, 0.643$\rightarrow$0.717 in SeedTTS test-en, 0.75$\rightarrow$0.774 in SeedTTS test-zh for SIM-O; +0.01 in LibriSpeech, +0.72 in SeedTTS test-en, +0.55 in SeedTTS test-zh for SMOS). When compared with human recordings, MaskGCT achieves human-level similarity across all three test sets (+0.017, -0.002, and +0.027 for SIM-O respectively in the three test sets, and +0.28, +0.32, and +0.25 for SMOS respectively in the three test sets). In robustness, MaskGCT likewise results nearly on par with ground truth (with 2.634, 2.623, 2.273 WER on LibriSpeech, SeedTTS test-en, and SeedTTS test-zh, respectively), exhibiting enhanced robustness compared to AR-based models and performing on par or better than NAR-based models such as VoiceBox and NaturalSpeech 3, without relying on phone-level duration predictions. In generation quality, MaskGCT achieves +0.10, +0.03, and +0.05 CMOS across the three test sets when compared with human recordings, indicating that MaskGCT attains human-level naturalness on these test sets. We also observe that MaskGCT exhibits excellent performance when using both ground truth total duration and predicted total duration, indicating the robustness of MaskGCT within a reasonable range of total speech duration and the capability of our total duration predictor to yield appropriate durations.
@@
AR + SoundStorm 0.672 3.267 0.998 4.20±${}_{\scriptscriptstyle\pm\text{0.17}}$${}_{\scriptscriptstyle\pm\text{0.20}}$
MaskGCT 0.687 2.634 0.886 4.27±${}_{\scriptscriptstyle\pm\text{0.14}}$ 0.10±${}_{\scriptscriptstyle\pm\text{0.16}}$
@@
AR + SoundStorm 0.683 2.846 0.323 4.03±${}_{\scriptscriptstyle\pm\text{0.23}}$0.05±${}_{\scriptscriptstyle\pm\text{0.22}}$
MaskGCT 0.717 2.623 0.188 4.24±${}_{\scriptscriptstyle\pm\text{0.12}}$ 0.03±${}_{\scriptscriptstyle\pm\text{0.14}}$
@@
AR + SoundStorm 0.747 3.865 0.238 3.78±${}_{\scriptscriptstyle\pm\text{0.23}}$0.32±${}_{\scriptscriptstyle\pm\text{0.19}}$
MaskGCT 0.774 2.273 0.106 4.09±${}_{\scriptscriptstyle\pm\text{0.12}}$ 0.05±${}_{\scriptscriptstyle\pm\text{0.17}}$
4.2.2 Autoregressive vs. Masked Generative Models
We compare MaskGCT to replacing T2S MaskGCT with an AR T2S model (which we call AR + SoundStorm). Table 3 shows the performance of these two models on all three test sets. MaskGCT demonstrates improved similarity, robustness, and CMOS (+0.12 on LibriSpeech test-clean, +0.08 on SeedTTS test-en, and +0.37 on SeedTTS test-zh) across all three test sets. We also conduct comparisons on more challenging hard cases (such as repeating words, and tongue twisters, which are often considered as samples where TTS systems are prone to hallucinations). MaskGCT exhibits a more pronounced robustness advantage in these scenarios. See details in Appendix J. In addition, compared to AR-based models, MaskGCT offers the capability to control the total duration of the generated speech, along with fewer inference steps, requiring only 25 to 50 steps for T2S models to achieve optimal results for speeches of any length. Conversely, the inference steps for AR-based models increase linearly with the length of the speech.
4.2.3 Duration Length Analysis
@@
Figure 3: WER vs. Total Duration Multiplier.
4.3 Speech Style Imitation
Zero-shot TTS endeavors to learn how to speak, including voice timbre and style, from prompt speech. Previous works utilized SIM-O to measure the similarity between generated speech and reference speech; however, SIM-O primarily assesses the similarity in voice timbre. In addition to evaluating the model’s zero-shot cloning ability through timbre similarity metrics, we also explored MaskGCT’s capability to clone overall style from two more expressive and stylized dimensions: accent and emotion. We randomly sampled a portion of data from the L2-ARCTIC Zhao et al. (2018) accent corpus and the ESD Zhou et al. (2021) emotion corpus to construct our accent and emotion evaluation datasets. Additionally, we introduce supplementary metrics to assess the model’s performance. For accent imitation, we employ SIM-Accent, to measure the similarity in accent between the generated speech and reference speech. The calculation process is analogous to SIM-O, but we utilize CommonAccent [^4] https://huggingface.co/Jzuluaga/accent-id-commonaccent_ecapa Xue et al. ; Qian et al. (2019) to derive the accent representation features of the speech. We also incorporate a subjective evaluation metric, Accent SMOS, which is similar to SMOS but focuses on accent rather than timbre. For emotion, we introduce Emotion SIM (with emotion2vec [^5] https://github.com/ddlBoJack/emotion2vec Ma et al. (2023) to extract features) and Emotion SMOS.
Our experiments demonstrate that MaskGCT exhibits powerful style cloning capabilities. For accent imitation, MaskGCT achieves the highest SIM-O of 0.717, close to the ground truth of 0.747. It also maintains a competitive WER of 6.382 and the best Accent SIM of 0.645. Additionally, MaskGCT leads in CMOS of 0.23, SMOS of 4.24, and Accent SMOS of 4.38. For emotion imitation, MaskGCT achieves the highest SIM-O of 0.600. It also attains a competitive WER of 12.502 and a strong Emotion SIM of 0.822. Furthermore, MaskGCT leads in all subjective metrics with CMOS of -0.31, SMOS of 4.07, and Emotion SMOS of 3.76, indicating natural and pleasant emotion imitation.
Table 4: Evaluation results for MaskGCT and the baseline methods on accent imitation.
System SIM-O $\uparrow$ WER $\downarrow$ Accent SIM $\uparrow$ CMOS $\uparrow$ SMOS $\uparrow$ Accent SMOS $\uparrow$
Accent Corpus L2-Arctit
Ground Truth 0.747 10.903 0.633 0.00 - -
VALL-E 0.403 10.721 0.485 -1.04±${}_{\scriptscriptstyle\pm\text{0.50}}$ 3.12±${}_{\scriptscriptstyle\pm\text{0.41}}$ 2.77±${}_{\scriptscriptstyle\pm\text{0.45}}$
CosyVoice 0.653 6.660 0.640 0.10±${}_{\scriptscriptstyle\pm\text{0.19}}$ 4.23±${}_{\scriptscriptstyle\pm\text{0.18}}$ 3.99±${}_{\scriptscriptstyle\pm\text{0.23}}$
VoiceBox 0.475 6.181 0.575 -0.55±${}_{\scriptscriptstyle\pm\text{0.22}}$ 3.93±${}_{\scriptscriptstyle\pm\text{0.25}}$ 3.49±${}_{\scriptscriptstyle\pm\text{0.29}}$
VoiceCraft 0.438 10.072 0.517 -0.39±${}_{\scriptscriptstyle\pm\text{0.22}}$ 3.51±${}_{\scriptscriptstyle\pm\text{0.33}}$ 3.29±${}_{\scriptscriptstyle\pm\text{0.28}}$
MaskGCT 0.717 6.382 0.645 0.23±${}_{\scriptscriptstyle\pm\text{0.17}}$ 4.24±${}_{\scriptscriptstyle\pm\text{0.16}}$ 4.38±${}_{\scriptscriptstyle\pm\text{0.25}}$
Table 5: Evaluation results for MaskGCT and the baseline methods on emotion imitation.
System SIM-O $\uparrow$ WER $\downarrow$ Emotion SIM $\uparrow$ CMOS $\uparrow$ SMOS $\uparrow$ Emotion SMOS $\uparrow$
Emotion Corpus ESD
Ground Truth 0.673 11.792 0.936 0.00 - -
VALL-E 0.396 15.731 0.735 -1.43±${}_{\scriptscriptstyle\pm\text{0.33}}$ 2.52±${}_{\scriptscriptstyle\pm\text{0.38}}$ 2.63±${}_{\scriptscriptstyle\pm\text{0.36}}$
CosyVoice 0.575 10.139 0.839 -0.45±${}_{\scriptscriptstyle\pm\text{0.18}}$ 3.98±${}_{\scriptscriptstyle\pm\text{0.19}}$ 3.66±${}_{\scriptscriptstyle\pm\text{0.19}}$
VoiceBox 0.451  12.647 0.811 -0.65±${}_{\scriptscriptstyle\pm\text{0.20}}$ 3.81±${}_{\scriptscriptstyle\pm\text{0.16}}$ 3.61±${}_{\scriptscriptstyle\pm\text{0.19}}$
VoiceCraft 0.345 16.042 0.788 -0.60±${}_{\scriptscriptstyle\pm\text{0.24}}$ 3.42±${}_{\scriptscriptstyle\pm\text{0.31}}$ 3.52±${}_{\scriptscriptstyle\pm\text{0.25}}$
MaskGCT 0.600 12.502 0.822 -0.31±${}_{\scriptscriptstyle\pm\text{0.17}}$ 4.07±${}_{\scriptscriptstyle\pm\text{0.16}}$ 3.76±${}_{\scriptscriptstyle\pm\text{0.25}}$
@@ 4.4 Ablation Study
Inference Timesteps. We explore the impact of inference steps of the T2S model on the results, ranging from 5 steps to 75 steps. Initially, SIM increases significantly and stabilizes after 25 steps. For test-zh, it rises from 0.761 at 5 steps to 0.771 at 75 steps, and for test-en, from 0.696 to 0.715. SIM peaks around 25 steps. WER improves more dramatically, especially up to 25 steps. For test-zh, it drops from 10.19 at 5 steps to 2.507 at 25 steps, and for test-en, from 8.096 to 2.346. Both SIM and WER show minimal changes beyond 25 steps. These findings suggest that SIM can be optimized with around 10 steps, while achieving the lowest WER requires approximately 25 steps. Beyond this, both metrics show minimal changes, indicating that further increases in steps do not yield substantial improvements. Therefore, for practical applications, 25 inference steps may be considered optimal for balancing SIM and WER, ensuring efficient and effective performance. See more details in Appendix A.2.
@@
Text Tokenizer. We compare two text tokenization methods: Grapheme-to-Phoneme (G2P) and Byte Pair Encoding (BPE). See more details in Appendix A.6.
@@ 5 Conclusion
In this paper, we present MaskGCT, a large-scale zero-shot TTS system that leverages fully non-autoregressive masked generative codec transformers while not requiring text-speech alignment supervision and phone-level duration prediction. MaskGCT achieves high-quality text-to-speech synthesis using text to predict semantic tokens extracted from a speech self-supervised learning (SSL) model, and then predicting acoustic tokens conditioned on these semantic tokens. Our experiments demonstrate that MaskGCT outperforms the state-of-the-art TTS system on speech quality, similarity, and intelligibility with scaled model size and training data, and MaskGCT can control the total duration of generated speech. We also explore the scalability of MaskGCT in tasks such as speech translation, voice conversion, emotion control, and speech content editing, demonstrating the potential of MaskGCT as a foundational model for speech generation.
@@ Appendix A Details of MaskGCT
A.1 Model Architecture
We use a Llama-style Touvron et al. [2023] Transformer architecture as the backbone of our model, incorporating gated linear units with GELU Hendrycks and Gimpel [2016] activation (SwiGLU), rotation position encoding Su et al. [2024], etc., but replacing causal attention with bidirectional attention. We also use adaptive RMSNorm Zhang and Sennrich [2019], which accepts the time step $t$ as the condition. Table 7 presents the key hyperparameters of the models.
@@ A.1 Model Architecture
A.2 Inference Steps for the T2S model
Figure 4 shows the relationship between inference steps and metrics SIM and WER for SeedTTS test-zh (left) and test-en (right). Initially, SIM increases significantly, stabilizing after 25 steps. For test-zh, SIM rises from 0.761 at 5 steps to 0.771 at 75 steps, and for test-en, from 0.696 to 0.715. SIM reaches high values with just 10 steps but peaks around 25 steps. WER improves more dramatically, especially up to 25 steps. For test-zh, WER drops from 10.19 at 5 steps to 2.507 at 25 steps, and for test-en, from 8.096 to 2.346. Both SIM and WER show minimal changes beyond 25 steps. These findings indicate that while SIM metrics can be sufficiently optimized with around 10 inference steps, achieving the lowest WER values requires approximately 25 inference steps. Beyond this threshold, both SIM and WER metrics exhibit minimal changes, implying that further increases in inference steps do not yield substantial improvements in these performance metrics. Therefore, for practical applications, 25 inference steps may be considered optimal for balancing SIM and WER, ensuring efficient and effective performance.
A.3 Inference Steps for the S2A model
The S2A model generates tokens layer by layer during inference. Since the acoustic codec follows an RVQ structure, we can view the S2A inference as a process from coarse to fine. We also use more iterations in the initial layers, as the first few layers carry more information. By default, we use inference steps of $[40,16,1,1,1,1,1,1,1,1,1,1]$ for each layer, however, we find that the S2A model can also perform well with fewer steps, such as $[10,1,1,1,1,1,1,1,1,1,1,1]$, with only a very slight performance loss.
@@ A.3 Inference Steps for the S2A model
A.4 Details of Semantic and Acoustic Codec
For semantic codec, we train a VQ-VAE model using the hidden features from the 17th layer of W2v-BERT 2.0, incorporating factorized codec Lezama et al. [2022] technology. The original hidden dimension of 1,024 is projected into a lower-dimensional space for quantization. The codebook size is set to 8,192, with a codebook dimension of 8. We employ only the $\mathcal{L}_{1}$ loss as the reconstruction target, optimizing the codebook with codebook loss and commitment loss. The input features are normalized to have a mean of 0 and a variance of 1, based on the statistics of the training dataset. The encoder and the decoder are each composed of 12 mirrored ConvNext blocks, featuring a kernel size of 7 and a hidden size of 384.
@@ A.4 Details of Semantic and Acoustic Codec
For acoustic codec, the basic architecture of the encoder follows Kumar et al. [2024] and the decoder follows Siuzdak [2023]. The Vocos-based decoder can model amplitude and phase, enabling waveform generation through inverse STFT transformation without requiring upsampling. The number of RVQ layers, codebook size, and codebook dimension are set to 12, 8,192, and 8, respectively. We utilize the multi-scale mel-reconstruction loss Kumar et al. [2024] $\mathcal{L}_{\text{rec}}$, for the adversarial loss $\mathcal{L}_{\text{adv}}$, we employ both the multi-period discriminator (MPD) and the multi-band multi-scale STFT discriminator, as proposed by Lee et al. [2022], Kumar et al. [2024]. Additionally, we incorporate the relative feature matching loss $\mathcal{L}_{\text{feat}}$. For codebook learning, we use the codebook loss $\mathcal{L}_{\text{codebook}}$ and the commitment loss $\mathcal{L}_{\text{commit}}$ from VQ-VAE. We set $\lambda_{\text{rec}}=10.0$, $\lambda_{\text{adv}}=2.0$, $\lambda_{\text{feat}}=2.0$, $\lambda_{\text{codebook}}=1.0$, $\lambda_{\text{commit}}=0.25$ as coefficients for balancing each loss terms. Figure 5 shows the overview of the semantic codec and acoustic codec, Table 9 presents the detailed model configurations of semantic codec and acoustic codec.
@@
A.5 Details of Duration Predictor